OpenStudy (anonymous):

60 degrees x (pie/180). I need a simple explanation, it's DMS conversion. And when i say simple, i mean SIMPLE

OpenStudy (jhannybean):

\[60^{\circ} \cdot \frac{\pi}{180^{\circ}} =\frac{60^{\circ}}{180^{\circ}}\cdot \pi =~?\]

OpenStudy (anonymous):

Okay, can you help me get the answer please lol

OpenStudy (anonymous):

cancel

OpenStudy (littlebird):

simplify that fraction